# High-dimensional structure underlying individual differences in naturalistic visual experience

### v1

@@ 2.3 Individual differences in cortical population codes and the interpretation of movies
To assess behavioral interpretations, we analyzed free recall transcripts collected after movie viewing. We used Google’s Universal Sentence Encoder (USE) to transform these transcripts into semantic embeddings that capture the conceptual content of each subject’s recall (Fig. 4a). For each movie, we constructed behavioral IDMs by computing pairwise correlations between subjects’ semantic embeddings, yielding a representation of the similarity structure in how individuals described and interpreted what they had seen. To test whether neural response patterns across different dimensional ranges predict behavioral similarities, we computed correlations between neural IDMs for each decade of latent dimensions and the behavioral IDMs.
@@
We found that neural similarity patterns predict behavioral similarity across dimensional ranges spanning two orders of magnitude in all regions examined (Fig. 4b, left panel). These brain-behavior correlations remained significant up to dimensions in the 11-100 range across all cortical regions ($p<0.05$, FDR-corrected). When controlling for coarse-scale effects, the correlation between neural and behavioral IDMs was attenuated across all regions but remained significant in the 11-100 dimensional range for the ventral, lateral, and posterior parietal-cingulate regions (Fig. 4b, right panel). In early visual cortex, the correlation between neural and behavioral IDMs was substantially reduced and no longer significant in the 11-100 range. Thus, while occipital regions contain coarse-grained information related to behavioral similarity, they lack fine-grained patterns that meaningfully correlate with individual differences in movie recall. In contrast, higher-level regions maintained behaviorally relevant information in fine-grained activity patterns after controlling for global response properties, suggesting that these subtle neural variations are important for capturing individual differences in stimulus interpretation.
@@
The persistence of brain-behavior correlations across multiple dimensional ranges demonstrates that behaviorally relevant individual differences extend well beyond the dominant dimensions that explain the majority of variance in neural responses. These findings reveal that each subject’s unique perspective on naturalistic content is reflected in high-dimensional population codes distributed over a hundred dimensions in each region of the cortical hierarchy, suggesting that fine-grained neural response patterns typically overlooked in conventional analyses may be crucial for capturing the rich diversity of individual visual experiences.
@@ 3 Discussion
By applying cross-decomposition to fMRI data collected during movie viewing, we characterized the complex representational structure underlying individual differences. Strikingly, we found a remarkable persistence of reliable individual differences across this expansive dimensional spectrum. Our analyses revealed that even dimensions with relatively low variance contain stable information about individual neural processing. The dimensional specificity observed in patterns of individual differences provides compelling evidence for multiple orthogonal axes of variation in how individuals process identical visual stimuli. This dimensional specificity persisted even after controlling for coarse-scale response properties captured by voxelwise ISC measures, confirming that fine-grained multivariate patterns contribute substantially to neural individuality beyond what can be explained by coarse-grained responses. Lastly, the preservation of significant correlations between neural and behavioral similarity matrices across dimensional ranges further demonstrated that these high-rank dimensions, despite explaining relatively little overall variance, encode functionally meaningful information that shapes individual interpretations of visual input.
@@
These findings bridge recent computational advances with fundamental questions about neural coding. Large-scale modeling efforts  (?, ?, ?, ?, ?, ?) and big-data approaches in neuroscience  (?, ?, ?) have enabled empirical estimates of latent dimensionality that are orders of magnitudes greater than previously possible. Recent work in systems neuroscience found that cortical representations are high-dimensional  (?, ?, ?, ?, ?), in line with theoretical perspectives suggesting that high-dimensional neural codes confer computational advantages for complex tasks  (?, ?, ?). Our work extends this framework to the neural basis of individual differences and demonstrates that neural population codes leverage their full high-dimensional capacity to organize shared and idiosyncratic features across multiple representational ranges. Critically, our findings connecting high-dimensional neural patterns to individual differences in movie recall suggest that subjective experience emerges not from dominant neural signals alone, but from the integration of information distributed across the entire variance spectrum.
@@
While previous work has shown that hyperalignment approaches improve detection of fine-grained individual differences by mapping responses to shared spaces  (?, ?, ?, ?), these previous studies did not examine the range of latent dimensions that contribute to these individual differences. Our spectral approach, which examines stimulus-related variance across the full spectrum of latent dimensions, reveals that individual differences manifest uniquely at each dimensional range—from dominant patterns in low-rank dimensions to subtle variations in high-rank dimensions. By explicitly modeling the distribution of variance across the full dimensional spectrum, our analysis reveals that high-rank dimensions encode qualitatively distinct and behaviorally significant information. Thus, our geometric perspective suggests focusing on the distribution and organization of variance across dimensions and provides a framework for characterizing the multiscale structure of neural individuality during naturalistic perception.
Several important questions remain. We demonstrated that different dimensional ranges capture distinct patterns of individual variability, but the specific representational content encoded within each range is left uncharacterized. Computational modeling approaches or targeted experimental manipulations could help interpret what cognitive features these dimensions represent—for instance, comparing neural patterns to representations from neural networks might reveal whether low-rank dimensions encode concrete, perceptual features and high-rank dimensions capture abstract, semantic information. Relatedly, it remains to be shown how high-dimensional representational structure transforms across the cortical hierarchy. We identified region-specific patterns of individual differences across dimensional ranges, but the evolution of these representations along the visual processing stream requires further investigation  (?). Lastly, whether individual differences in variance distributions can be linked to trait-like phenotypes is an open question. Investigating the full dimensional structure of neural responses could identify subtle perturbations in how individuals with perceptual or cognitive disorders organize information across representational scales. For example, atypical sensory processing in neuropsychiatric conditions may manifest as altered distributions of information across neural dimensions, revealing neurobiological signatures undetectable through conventional measures of regional activation or connectivity  (?, ?, ?, ?, ?). Future work should investigate these questions to develop a more complete understanding of how neural geometry shapes individual visual experience.
@@
Methods
3.1 Naturalistic fMRI dataset
@@
We analyzed fMRI data from 42 subjects viewing four audiovisual movies  (?). We excluded one subject whose recall data were missing from the original dataset (n=43). Movie durations ranged from 7:27 to 12:27 minutes and varied in content and cinematic style, including social and affective elements and screen cuts. Three out of four movies depicted scenarios with social interactions (versus one with purely mechanical actions), and two contained screen cuts (versus continuous camera panning).
fMRI preprocessing
@@
Functional images (TR=1 s) were preprocessed using fMRIPrep  (?), including motion correction, co-registration, and normalization to MNI space at 2mm isotropic resolution. Preprocessed data were trimmed to movie segments and high-pass filtered (64 s cutoff) using FSL  (?).
@@
We defined four anatomical regions of interest (ROIs) spanning the visual cortical hierarchy using the Automatic Anatomical Labeling atlas (AAL3)  (?): the early visual regions (including calcarine, superior, middle, and inferior occipital cortices), the ventral visual stream (including lingual, fusiform, inferior temporal, and parahippocampal cortices), and the lateral visual stream (including middle and superior temporal cortices). We also created an ROI by combining higher-level regions in the posterior part of the cortex (including supramarginal and angular gyri, precuneus, and posterior cingulate cortex) that we collectively refer to as posterior parietal-cingulate (”parietal-cingulate”). We grouped these high-level posterior regions to maintain comparable voxel counts across ROIs while combining regions that support similar semantic and multimodal integration functions. The covariance spectra patterns show similar high-dimensional structure when these posterior regions are analyzed separately. ROI masks were applied to each subject’s preprocessed movie fMRI responses to extract the corresponding voxel timeseries matrices. All regions included both left and right hemispheres.
3.2 Cross-decomposition
Estimating latent dimension
We used cross-decomposition to estimate the dimensionality of the latent neural subspace shared by pairs of individuals. For each subject pair, we analyzed their data matrices $X,Y\in\mathbb{R}^{t\times v}$ representing $t$ timepoints and $v$ common voxels from subjects $i$ and $j$, respectively. We computed the training set cross-covariance eigenvalues and eigenvectors using singular vector decomposition:

### v2

@@ 2.3 Individual differences in cortical population codes and the interpretation of movies
To assess behavioral interpretations, we analyzed free recall transcripts collected after movie viewing. We used Google’s Universal Sentence Encoder (USE) to transform these transcripts into semantic embeddings that capture the conceptual content of each subject’s recall (Fig. 4A). For each movie, we constructed behavioral IDMs by computing pairwise correlations between subjects’ semantic embeddings, yielding a representation of the similarity structure in how individuals described and interpreted what they had seen. Notably, the average between-subject correlation for the same movie was 0.59 ± 0.10, compared to the average within-subject correlation across different movies of 0.47 ± 0.09. This shows that the divergence across subjects for the same movie is nearly as strong as the divergence across different movies within each subject, confirming substantial inter-individual variability in movie interpretation. To test whether neural response patterns across different dimensional ranges predict behavioral similarities, we computed correlations between neural IDMs for each decade of latent dimensions and the behavioral IDMs. Because neural-behavioral correlations are more subtle than IDM reliability correlations, we used coarser dimensional binning to improve statistical power while still capturing distinctions across the dimensional spectrum.
@@
We found that neural similarity patterns predict behavioral similarity across dimensional ranges spanning two orders of magnitude in all regions examined (Fig. 4B, left panel). These brain-behavior correlations remained significant up to dimensions in the 11-100 range across all cortical regions ($p<0.05$, FDR-corrected). When controlling for coarse-scale effects, the correlation between neural and behavioral IDMs was attenuated across all regions but remained significant in the 11-100 dimensional range for the ventral, lateral, and posterior parietal-cingulate regions (Fig. 4B, right panel). In early visual cortex, the correlation between neural and behavioral IDMs was substantially reduced and no longer significant. Thus, while occipital regions contain coarse-grained information related to behavioral similarity, they lack fine-grained patterns that meaningfully correlate with individual differences in movie recall. In contrast, higher-level regions maintained behaviorally relevant information in fine-grained activity patterns after controlling for global response properties, suggesting that these subtle neural variations are important for capturing individual differences in stimulus interpretation. Partial correlation analyses controlling for other dimensional ranges further confirmed that different ranges of dimensions provided independent contributions to this relationship (Fig. S5A).
@@
Together, the persistence of brain-behavior correlations across multiple dimensional ranges demonstrates that behaviorally relevant individual differences extend well beyond the dominant dimensions that explain the majority of variance in neural responses. These findings reveal that each subject’s unique perspective on naturalistic content is reflected in high-dimensional population codes distributed over a hundred dimensions in each region of the cortical hierarchy, suggesting that fine-grained neural response patterns typically overlooked in conventional analyses may be crucial for capturing the rich diversity of individual visual experiences.
2.4 Neural alignment across high-dimensional codes predicts concrete narrative descriptions
Having established that high-dimensional neural patterns predict individual differences in movie recall, we next sought to understand what interpretable aspects of behavior might relate to these individual differences in neural representations. Inspection of subjects’ verbal recall transcripts revealed an intriguing pattern in descriptive style. While some subjects primarily focused on concrete, perceptual details of the movies such as specific actions, objects, and events, others reported more abstract, interpretive aspects such as themes, valence, and personal reactions (see examples in Fig. 5A). This observation led us to hypothesize that the degree of concreteness in subjects’ descriptions might be systematically related to their neural similarity.
We reasoned that individuals who encode and recall concrete details would show greater neural alignment with one another because concrete details reflect more consistent, stimulus-driven processing of the shared sensory input. In contrast, abstract and interpretive processing may be more idiosyncratic, leading to greater inter-individual variability in neural responses. To test this hypothesis, we quantified the concreteness of each subject’s recall transcripts using a large-scale concreteness ratings database [brysbaert2014concreteness] and correlated it with their average neural covariance with all other subjects across different dimensional ranges (see Methods). We found that individuals who produced more concrete narrative descriptions showed significantly greater neural similarity to others across multiple dimensional ranges (Fig. 5B).
We then examined whether individuals who recall more concretely had higher shared dimensionality. We computed the number of significant shared dimensions between each subject and all others and correlated this with each subject’s concreteness rating. Individuals with more concrete narrative styles tended to show higher shared neural dimensionality across all cortical regions examined (Fig. 5C). These findings suggest that an important factor underlying individual differences in both neural representations and behavioral recall is the degree to which subjects encode the rich concrete details of their experiences versus more abstract, interpretive aspects.
@@ 2.4 Neural alignment across high-dimensional codes predicts concrete narrative descriptions
Figure 5: Concrete narrative descriptions predict greater neural alignment across high-dimensional codes. (A) Sample excerpts from two subjects illustrating the range of narrative concreteness, from highly concrete descriptions emphasizing perceptible events and objects (top) to more abstract, thematic accounts (bottom). (B) Individuals who produced more concrete narrative descriptions showed greater neural similarity to others across multiple dimensional ranges. Bar plots show correlations between individual concreteness scores and mean intersubject neural covariance across dimensional ranges. Left: original covariance; right: after removing coarse-scale effects. Dots above each bar indicate statistical significance ($p<0.05$, FDR-corrected). Small triradiate symbols represent results for individual movies without averaging. (C) Subject-level correlations between narrative concreteness and the number of shared neural dimensions with above-chance inter-subject covariance. Individuals with more concrete narrative styles tended to have higher shared neural dimensionality across all cortical regions (Spearman $\rho=0.254-0.365$). Each point represents one subject.
@@ 3 Discussion
By applying cross-decomposition to fMRI data collected during movie viewing, we characterized the complex representational structure underlying individual differences. Strikingly, we found a remarkable persistence of reliable individual differences across this expansive dimensional spectrum. Our analyses revealed that even dimensions with relatively low variance contain stable information about individual neural processing. The dimensional specificity observed in patterns of individual differences provides compelling evidence for multiple orthogonal axes of variation in how individuals process identical visual stimuli. This dimensional specificity persisted even after controlling for coarse-scale response properties captured by voxelwise ISC measures, confirming that fine-grained multivariate patterns contribute substantially to neural individuality beyond what can be explained by coarse-grained responses. Lastly, the preservation of significant correlations between neural and behavioral similarity matrices across dimensional ranges further demonstrated that these high-rank dimensions, despite explaining relatively little overall variance, encode functionally meaningful information that shapes individual interpretations of visual input. Notably, individuals who recall more concrete details exhibit higher-dimensional shared representations.
@@
These findings bridge recent computational advances with fundamental questions about neural coding. Large-scale modeling efforts [krizhevsky2017imagenet, he2016deep, dosovitskiy2020image, vaswani2017attention, brown2020language, kaplan2020scaling] and big-data approaches in neuroscience [allen2022massive, hebart2023things, manley2024simultaneous] have enabled empirical estimates of latent dimensionality that are orders of magnitudes greater than previously possible. Recent work in systems neuroscience found that cortical representations are high-dimensional [stringer2019high, contier2024distributed, elmoznino2024high, manley2024simultaneous, gauthaman2024], in line with theoretical perspectives suggesting that high-dimensional neural codes confer computational advantages for complex tasks [FUSI201666, sorscher2022neural, elmoznino2024high]. Our work extends this framework to the neural basis of individual differences and, contrasting with earlier findings of organization along a small number of dimensions based on leading high-variance components [huth2012continuous, tarhan2020sociality, khosla2022highly], demonstrates that neural population codes leverage their high-dimensional capacity to organize shared and idiosyncratic features across multiple representational ranges. Critically, our findings connecting high-dimensional neural patterns to individual differences in movie recall suggest that subjective experience emerges not from dominant neural signals alone, but from the integration of information distributed across the entire variance spectrum.
@@
Our findings that reliable neural representations span at least a hundred dimensions within individual regions are particularly striking given that this high-dimensional structure extends throughout the visual cortical hierarchy, including higher-level semantic regions in posterior parietal and cingulate cortex. The richness of naturalistic movie stimuli, combined with the cross-decomposition method, enables us to detect this expansive representational structure across both early sensory and higher-order integrative regions. These results are consistent with, and extend, prior functional alignment methods that have identified shared representations spanning hundreds of latent dimensions across the whole cortex. [guntupalli2016model] However, while previous work has used hyperalignment to improve the detection of fine-grained individual differences [haxby2011common, guntupalli2016model, feilong2018reliable, Guntupalli_2020, Feilong_Guntupalli_Haxby_2021], these previous studies did not examine the range of latent dimensions that contribute to these individual differences. Our spectral approach, which examines stimulus-related variance across the full spectrum of latent dimensions, reveals that individual differences manifest uniquely at each dimensional range—from dominant patterns in low-rank dimensions to subtle variations in high-rank dimensions. By explicitly modeling the distribution of variance across the full dimensional spectrum, our analysis reveals that high-rank dimensions encode qualitatively distinct and behaviorally significant information. Thus, our work suggests focusing on the distribution and organization of variance across dimensions and provides a framework for characterizing the multiscale structure of neural individuality during naturalistic perception.
Our findings suggest several important directions for future work. First, we found that reliability of neural individual differences based on higher ranks are relatively lower than that based on lower ranks. While several observations indicate lower effective signal-to-noise ratio at higher ranks, it remains to be investigated whether the lower reliabilities reflect decreased signal quality or genuinely smaller individual differences at finer representational scales. Second, our temporal autocorrelation analysis suggests that higher-rank dimensions operate at finer temporal scales, raising the question of whether higher ranks might encode stimulus information that varies more rapidly. Third, an exciting direction for future work will be to investigate whether specific dimensions or ranges of dimensions can be mapped onto interpretable features of the movie stimuli, or whether they have complex mixed selectivities that cannot be easily interpreted. Lastly, our current approach uses anatomically-defined ROIs and extracts shared components separately for each subject pair. Future extensions could use functionally-defined ROIs based on individualized cortical topographies and template-based alignment methods that enable comparison of the same components across all subject pairs, as implemented in the Individualized Neural Tuning model [feilong2023individualized].
@@
STAR Methods
3.1 Key resources table
REAGENT or RESOURCE SOURCE IDENTIFIER
Deposited Data
Naturalistic movie fMRI dataset Sava-Segal et al. [savasegal] OpenNeuro: ds004516
Software and Algorithms
fMRIPrep Esteban et al. [Esteban2019] RRID:SCR_016216
FSL Smith et al. [smith2004advances] RRID:SCR_002823
Python Python Software Foundation https://www.python.org
Google Universal Sentence Encoder Cer et al. [cer2018universal] https://tfhub.dev/google/universal-sentence-encoder/4
custom code This paper https://github.com/kelseyhan-jhu/idiosyncratic-neural-geometry
Other
AAL3 Atlas Rolls et al. [ROLLS2020116189] https://www.gin.cnrs.fr/en/tools/aal/
3.2 Experimental model and study participant details
3.2.1 fMRI participants
We analyzed publicly available fMRI data from Sava-Segal et al. [savasegal]. From 48 initially recruited participants (all native English speakers and right-handed; 27 female, median age 24.5 years, range 19-64 years), the original study retained 43 participants with complete fMRI data for all four movies. We further excluded 1 participant with missing behavioral recall data, resulting in a final sample of 42 participants (all native English speakers and right-handed, 24 female, median age 24.5 years, range 19-64 years). All participants provided informed written consent under NIH Institutional Review Board approval.
3.3 Method details
3.3.1 Movie stimuli
@@
Participants viewed four naturalistic audiovisual movie clips (duration 7:27-12:27 min). Movies varied in content and cinematic style, including social and affective elements and screen cuts. Three out of four movies depicted scenarios with social interactions (versus one with purely mechanical actions), and two contained screen cuts (versus continuous camera panning). Detailed stimulus descriptions and links are available in Sava-Segal et al. [savasegal].
Behavioral recall data
Following each movie, participants provided free verbal recall for three minutes describing characters, events, and opinions.
Narrative concreteness
To quantify how concretely versus abstractly subjects described their experiences, we computed concreteness scores for each subject’s recall transcripts using a large-scale database [brysbaert2014concreteness]. This database provides 5-point scale ratings (1=abstract, 5=concrete) for 37,058 English words collected via Amazon Mechanical Turk. For each subject’s transcript, we computed the mean concreteness rating across all words present in the database (mean coverage: 90.9% of words per transcript, range: 85.4%-96.3%). We then averaged concreteness scores across all four movies for each subject to obtain a subject-level concreteness measure.
3.3.2 fMRI experiment
fMRI data acquisition
Functional MRI data were collected at the National Institutes of Health using a 3T Siemens Prisma scanner with a 64-channel head coil. Movies were presented in pseudorandomized order (counterbalanced across participants) via rear-projection screen viewed through an angled mirror using PsychoPy [peirce2019]. Functional images were acquired using T2*-weighted multiband, multi-echo echo-planar imaging with TR = 1,000 ms, echo times = [13.6, 31.86, 50.12 ms], flip angle = 60°, 3.0 mm isotropic voxels, 52 slices (whole-brain coverage), multiband acceleration factor = 4. Anatomical images were acquired using T1-weighted MPRAGE with TR = 2,530 ms, TE = 3.30 ms, flip angle = 7°, 1.0 mm isotropic resolution. See Sava-Segal et al. [savasegal] for complete acquisition details.
fMRIPrep
Results included in this manuscript come from preprocessing performed using fMRIPrep 21.0.2 [fmriprep1, fmriprep2], which is based on Nipype 1.6.1 [nipype1, nipype2].
The next two sections (anatomical data and functional data) of boilerplate text were automatically generated by fMRIPrep with the express intention that users should copy and paste this text into their manuscripts unchanged. It is released under the CC0 license.
Anatomical data
A total of 1 T1-weighted (T1w) images were found within the input BIDS dataset. The T1-weighted (T1w) image was corrected for intensity non-uniformity (INU) with N4BiasFieldCorrection [n4], distributed with ANTs 2.3.3 [ants], and used as T1w-reference hroughout the workflow. The T1w-reference was then skull-stripped with a Nipype implementation of the antsBrainExtraction.sh workflow (from ANTs), using OASIS30ANTs as target template. Brain tissue segmentation of cerebrospinal fluid (CSF), white-matter (WM) and gray-matter (GM) was performed on the brain-extracted T1w using fast (FSL 6.0.5.1:57b01774 [fsl_fast]). Volume-based spatial normalization to one standard space (MNI152NLin2009cAsym) was performed through nonlinear registration with antsRegistration (ANTs 2.3.3), using brain-extracted versions of both T1w reference and the T1w template. The following template was selected for spatial normalization: ICBM 152 Nonlinear Asymmetrical template version 2009c [mni152nlin2009casym] (TemplateFlow ID: MNI152NLin2009cAsym).
Functional data
For each of the 1 BOLD runs found per subject (across all tasks and sessions), the following preprocessing was performed. First, a reference volume and its skull-stripped version were generated from the shortest echo of the BOLD run using a custom methodology of fMRIPrep. Head-motion parameters with respect to the BOLD reference (transformation matrices, and six corresponding rotation and translation parameters) are estimated before any spatiotemporal filtering using mcflirt (FSL 6.0.5.1:57b01774) [mcflirt]. BOLD runs were slice-time corrected to 0.452s (0.5 of slice acquisition range 0s-0.905s) using 3dTshift from AFNI [afni]. The BOLD time-series (including slice-timing correction when applied) were resampled onto their original, native space by applying the transforms to correct for head-motion. These resampled BOLD time-series will be referred to as preprocessed BOLD in original space, or just preprocessed BOLD. A T2* map was estimated from the preprocessed EPI echoes, by voxel-wise fitting the maximal number of echoes with reliable signal in that voxel to a monoexponential signal decay model with nonlinear regression. The T2*/S0 estimates from a log-linear regression fit were used for initial values. The calculated T2* map was then used to optimally combine preprocessed BOLD across echoes following the method described in Posse et al. [posse_t2s]. The optimally combined time series was carried forward as the preprocessed BOLD. The BOLD reference was then co-registered to the T1w reference using mri_coreg (FreeSurfer) followed by flirt (FSL 6.0.5.1:57b01774) [flirt], with the boundary-based registration cost-function [bbr]. Co-registration was configured with six degrees of freedom. First, a reference volume and its skull-stripped version were generated using a custom methodology of fMRIPrep. Several confounding time-series were calculated based on the preprocessed BOLD: framewise displacement (FD), DVARS and three region-wise global signals. FD was computed using two formulations following Power (absolute sum of relative motions [power_fd_dvars] and Jenkinson (relative root mean square displacement between affines [mcflirt]. FD and DVARS are calculated for each functional run, both using their implementations in Nipype following the definitions by Power et al. [power_fd_dvars]). The three global signals are extracted within the CSF, the WM, and the whole-brain masks. Additionally, a set of physiological regressors were extracted to allow for component-based noise correction CompCor [compcor]. Principal components are estimated after high-pass filtering the preprocessed BOLD time-series (using a discrete cosine filter with 128s cut-off) for the two CompCor variants: temporal (tCompCor) and anatomical (aCompCor). tCompCor components are then calculated from the top 2% variable voxels within the brain mask. For aCompCor, three probabilistic masks (CSF, WM and combined CSF+WM) are generated in anatomical space. The implementation differs from that of Behzadi et al. in that instead of eroding the masks by 2 pixels on BOLD space, the aCompCor masks are subtracted a mask of pixels that likely contain a volume fraction of GM. This mask is obtained by thresholding the corresponding partial volume map at 0.05, and it ensures components are not extracted from voxels containing a minimal fraction of GM. Finally, these masks are resampled into BOLD space and binarized by thresholding at 0.99 (as in the original implementation). Components are also calculated separately within the WM and CSF masks. For each CompCor decomposition, the k components with the largest singular values are retained, such that the retained components’ time series are sufficient to explain 50 percent of variance across the nuisance mask (CSF, WM, combined, or temporal). The remaining components are dropped from consideration. The head-motion estimates calculated in the correction step were also placed within the corresponding confounds file. The confound time series derived from head motion estimates and global signals were expanded with the inclusion of temporal derivatives and quadratic terms for each [confounds_satterthwaite_2013]. Frames that exceeded a threshold of 0.5 mm FD or 1.5 standardised DVARS were annotated as motion outliers. The BOLD time-series were resampled into standard space, generating a preprocessed BOLD run in MNI152NLin2009cAsym space. First, a reference volume and its skull-stripped version were generated using a custom methodology of fMRIPrep. All resamplings can be performed with a single interpolation step by composing all the pertinent transformations (i.e. head-motion transform matrices, susceptibility distortion correction when available, and co-registrations to anatomical and output spaces). Gridded (volumetric) resamplings were performed using antsApplyTransforms (ANTs), configured with Lanczos interpolation to minimize the smoothing effects of other kernels [lanczos]. Non-gridded (surface) resamplings were performed using mri_vol2surf (FreeSurfer).
Many internal operations of fMRIPrep use Nilearn 0.8.1 [nilearn], mostly within the functional processing workflow. For more details of the pipeline, see the section corresponding to workflows in fMRIPrep’s documentation.
Post-processing
@@
Preprocessed data were trimmed to movie segments and high-pass filtered (64 s cutoff) using FSL [smith2004advances].
@@
We defined four anatomical regions of interest (ROIs) spanning the visual cortical hierarchy using the Automatic Anatomical Labeling atlas (AAL3) [ROLLS2020116189]: the early visual regions (including calcarine, superior, middle, and inferior occipital cortices), the ventral visual stream (including lingual, fusiform, inferior temporal, and parahippocampal cortices), and the lateral visual stream (including middle and superior temporal cortices). We also created an ROI by combining higher-level regions in the posterior part of the cortex (including supramarginal and angular gyri, precuneus, and posterior cingulate cortex) that we collectively refer to as posterior parietal-cingulate (”parietal-cingulate”). We grouped these high-level posterior regions to maintain comparable voxel counts across ROIs while combining regions that support similar semantic and multimodal integration functions. The covariance spectra patterns show similar high-dimensional structure when these posterior regions are analyzed separately. Maintaining similar ROI sizes was important because the number of detectable dimensions scales with the number of voxels. ROI masks were applied to each subject’s preprocessed movie fMRI responses to extract the corresponding voxel timeseries matrices. All regions included both left and right hemispheres.
For whole-brain analysis (Supplementary Fig. S4), we defined 170 bilateral ROIs corresponding to all AAL3 parcels (Rolls et al., 2020). We excluded 11 ROIs where more than 10% of subject pairs had insufficient overlapping voxels and thus insufficient rank to compute cross-covariance. The excluded regions were primarily subcortical structures with limited coverage in the functional data (anterior cingulate, thalamus, reuniens thalamic nucleus, medial geniculate nucleus, ventral tegmental area, and median raphe nucleus).
3.4 Quantification and statistical analysis
3.4.1 Estimating latent dimension
Cross-decomposition
We used cross-decomposition to estimate the dimensionality of the latent neural subspace shared by pairs of individuals. For each subject pair, we analyzed their data matrices $X,Y\in\mathbb{R}^{t\times v}$ representing $t$ timepoints and $v$ common voxels from subjects $i$ and $j$, respectively. Here we used common voxels because the data were in MNI space, but this method does not require the two data matrices to be in a common space. Voxel timecourses were z-scored for each subject and each movie before computing cross-subject covariance. We computed the training set cross-covariance eigenvalues and eigenvectors using singular vector decomposition: